OpenStudy (anonymous):

Solve the system, using substitution. y = 1 - 5x and 5x + y = 8 A. no solution B. infinitely many solutions

jimthompson5910 (jim_thompson5910):

start with the second equation and plug in y = 1-5x 5x + y = 8 5x + (1-5x) = 8 5x - 5x + 1 = 8 I'll let you finish